# Ghost projection via focal-field diffraction catastrophes

### v1

@@ IV.1 Truncated Zernike basis for phase modulations
The beam in Fig. 1(a) is taken to be transversely confined by a circular aperture. Zernike polynomials [11, 24, 25, 26, 27] form a complete orthogonal set over such a region. They therefore provide a convenient basis for describing the aperture-plane phase modulations that can be used to generate diffraction catastrophes for the purposes of ghost projection. Working in plane polar coordinates $(\rho,\theta)$, denote the first 21 Zernike polynomials by
@@
To bound the phase modulations, each $c^{k}_{j}$ is randomly chosen from the interval $[-\tau,\tau]$ for some value $\tau$. In order for the focal-field intensity to contain a caustic network, the value for $\tau$ needs to be sufficiently large. We set the amplitude of the modulated wavefunction to be uniformly 1 within the unit circle and zero elsewhere, thereby simulating a circular aperture that is uniformly illuminated by normally-incident strictly monochromatic plane waves, so
@@
For the results in this paper, a randomly-generated set of $\mathcal{N}=40,000$ illuminating patterns were drawn from to construct $100\times 100$ pixel desired (or target) images. Note that the number of patterns, $\mathcal{N}$, is greater than the number of pixels in order to increase the expected number of dimensions spanned by the $I_{k}$, as being randomly generated, there is no guarantee that the $I_{k}$ are linearly independent. [^7] If sufficiently many random patterns are chosen, the resulting ensemble approximately non-negatively spans the space, i.e. every target image can be well-approximated by a linear combination of patterns with non-negative coefficients. Probabilistically, we actually would expect any randomly chosen 10,000 vectors to be linearly independent, but they almost certainly would not non-negatively span the space associated with $100\times 100$ pixel images, which is the key aspect in the present context. Additionally, as an area for further exploration, $\mathcal{N}$ being far greater than the number of pixels means most vectors will be able to be constructed up to a pedestal via many different linear combinations of the illuminating patterns. The pedestal profile is expected to be largely the same for different linear combinations supposing the number of vectors in each linear combination is sufficiently large, but the pedestal’s magnitude will depend on the linear combination, so it makes sense to talk about choosing the linear combination which maximizes the signal to pedestal ratio (which leads to improved signal to noise ratio). A mechanism for this has not been explored in this paper. For $k\in\{1,\cdots,\mathcal{N}\}$ and $j\in\{2,\cdots,21\}$, each of the $c^{k}_{j}$ was uniformly randomly selected from $[-\tau,\tau]$. It was found that the size of $\tau$ is approximately proportional to the average size of the corresponding focal-field illuminating pattern, which is natural on account of the fact that the mean aperture-plane phase gradient is linearly proportional to $\tau$. For this paper we choose $\tau=9$, as for $100\times 100$ pixel images, this makes the patterns a suitable size for the target images to be synthesized.
@@ IV.2 Focusing the phase modulations via the Fourier transform
A coherent radiation distribution, when propagated infinitely far, becomes the Fourier transform of the initial distribution. A focusing lens enables this to occur at a finite distance, and the Fourier transform appears at the focal plane [11]. In particular, once the modulated wavefunction, $\psi_{k}$, is focused via a lens, the intensity distribution, $I_{k}$, in the focal plane is given by
@@
Here, $\mathcal{F}$ denotes Fourier transformation with respect to transverse Cartesian coordinates $\rho\cos\theta$ and $\rho\sin\theta$, and $A(\rho)$ is an aperture function in scaled (dimensionless) spatial coordinates, defined to be unity for $\rho\leq 1$ and zero elsewhere. The resulting $I_{k}$ are diffraction catastrophes [12, 13, 10]. Here $\{I_{k}\}_{k\in\{1,\cdots,\mathcal{N}\}}$ serves as the caustic-laden “catastrophe basis” set of illuminating patterns to be used for ghost projection. The process described above is depicted in Fig. 1(a) and some example patterns are shown in Fig. 1(b). For purposes of simulation, it is helpful to rasterize the wavefunctions, treating $A(\rho)\exp[i\phi_{k}(\rho,\theta)]$ as a vector with a complex-valued entry for each pixel. This allows the Fourier transform step to be performed computationally as a 2D fast Fourier transform (FFT) [28]. To avoid wrapping of the FFT at the image boundaries, the phase function must be zero-padded, i.e., the phase function is given a sufficiently broad border of value zero.
@@ IV.3 Simulating Poisson noise in the focal fields
The number of photons (or electrons, neutrons, etc.) incident on a pixel of the projection plane obeys a Poisson probability distribution [20, 29]. For simplicity, assume the projecting beam is composed of photons. For the $j$th pixel of the ghost projection, let $\Lambda_{j}=c\sum_{k}w_{k}I_{k}$ represent the expected number of photons incident on that pixel for some scaling factor $c$. When measured, the number of photons incident at each pixel, $j$, will be Poisson distributed, with mean $\Lambda_{j}$ and standard deviation $\sqrt{\Lambda_{j}}$. To quantify noise levels, a pixel-wise signal-to-noise ratio (SNRj) is introduced, given by
$\text{SNR}_{j}=\frac{|\Lambda_{j}-P|}{\sqrt{\Lambda_{j}}}$ (17)
@@
$\text{SNR}=\sqrt{\sum_{j=0}^{N^{2}}\frac{\text{SNR}_{j}^{2}}{N^{2}}}=\frac{1}{N}\left\lVert\frac{\Lambda-P}{\sqrt{\Lambda}}\right\rVert_{2}.$ (18)
@@ V Results and Analysis
In this section we first compare ghost projection performance using both (i) the analytical dot product method, and (ii) the iterative non-negative least-squares method, to determine pattern weights in Sec. V.1. Unlike a random speckle pattern basis that produces a flat (or uniform) pedestal in target-image synthesis, we observe that a diffraction-catastrophe basis produces a non-uniform pedestal. This observation is presented in Sec. V.2. We then present a method to modify the target image to compensate for this non-uniform pedestal in Sec. V.3. Finally, we explore the effects of Poisson noise on target image quality in Sec. V.4.
@@
V.1 Synthesis of unmodified target images
Figure 3(a) shows an example of a phase function, $\phi_{k}$, that is generated randomly using Zernike polynomials from $\mathcal{Z}$, with the corresponding diffraction catastrophe, $I_{k}$, after focusing shown in Figs. 3(b-c). The spatial scale is in arbitrary units, for all panels. The caustic diffraction catastrophes [10] are evident as the bright lines in the intensity images. While these lines would correspond to nonphysical lines of infinite intensity in the geometric-optics limit of zero wavelength, the finite wavelengths softens them to peaked intensities decorated with fine grid-like structures associated with a lattice of alternating phase vortices and antivortices [13].
@@ V.1 Synthesis of unmodified target images
Using a set of $\mathcal{N}=40,000$ randomly generated aberrated focal-field diffraction-catastrophe patterns as a random basis, the simulated ghost projections in Fig. 4 are produced. The target image, here, is a binary capital letter “R” in Fig. 4(a). Two different ghost-projection algorithms are explored below, namely dot product ghost projection [1] and NNLS ghost projection [2, 3, 4], as described in the following paragraphs. The dot product reconstruction in Fig. 4(b) has comparatively far lower contrast, with the white section of the target image sitting only approximately 5% above the local pedestal, whereas in the NNLS construction of Fig. 4(c), the white section sits approximately 34% above the local pedestal. This greater contrast is particularly valuable when noise is present (as explored further in Sec. IV.3).
@@
Figure 4: (a) A $100\times 100$ pixel image of the letter “R” used as a target image, (b) a simulated ghost-projection reconstruction of the target image weighting the catastrophes by their dot product with the target image, (c) a reconstruction using the non-negative least squares algorithm to generate weights. The reconstructions have been normalized, so that the minimum value is rendered black and the maximum value is rendered white. (d) There is low correlation between the two methods for generating the weights, $w_{k}$, with the Pearson correlation coefficient $r$ being 0.05.
@@
The “dot product” ghost-projection construction, which follows Ref. [1] by using weights defined in Eq. (10), discards illuminating patterns that have a below-average dot product with the target image. If one thinks of the random illumination patterns as randomly directed vectors in a function space, this method for ghost projection has the simple geometric interpretation that any function-space vector may be synthesized by taking a sheaf of random vectors and retaining only those having a positive projection along the direction of the required vector; the sum of these retained vectors is then approximately parallel to the desired vector (target ghost-projection image) (see Fig. 11 in Ref. [1]). Illuminating patterns with lower dot products with the desired image increase the pedestal while having a reduced contribution towards improving the ghost-projection image. Hence the weights of the included illuminating patterns are the dot product between the image and illuminating pattern vector, minus the average dot product. While conceptually clear, the very low contrast of the “dot product” ghost-projection method (see Fig. 4(b)) argues against its utility in practice.
@@
Further comparison of the two ghost-projection approaches may be obtained by plotting each weight, $w_{k}$, obtained using the NNLS approach, versus the corresponding weight obtained via the dot product approach. The resulting scatter plot in Fig. 4(d) reveals a notable disparity between the two reconstruction algorithms, with no significant correlation between the weights $w_{k}$ generated by each procedure. While Pearson’s correlation coefficient ($r=0.05$, stated in Fig. 4(d)) is positive, the support for a linear relationship between the dot product and the NNLS weights is weak, so the dot product cannot easily be used to filter illuminating patterns before the NNLS algorithm is applied. It is worth noting that the dot products are not symmetrically distributed around their mean and are positively skewed. Additionally, the NNLS algorithm only assigns 158 illuminating patterns a weight that is over 1% of the maximum weight (only 0.395% of the 40,000 illuminating patterns), with 62 of these having a dot product below the mean correlation (red) and 96 having a dot product above the mean correlation (blue). This small selection of contributing illuminating patterns is the reason that the NNLS image is noticeably more grainy than the dot product reconstruction, without a large quantity of similarly-weighted illuminating patterns to smooth the image.
@@ V.2 Non-uniform pedestal
Another significant aspect of the images formed in Fig. 4 is the additional background signal, or pedestal; in particular, that it is not spatially uniform. For the case of the dot product construction in Fig. 4(b), the pedestal is approximately the average illuminating pattern (see Fig. 5), while the NNLS construction in Fig. 4(c) has a flatter pedestal of similar profile (as there are far fewer contributing illuminating patterns, the resulting pedestal when using NNLS is more dependent on the target image).
Figure 5: The average diffraction catastrophe intensity function. The 2D image above has not been normalized. As may be expected from the central limit theorem, the average mask is well approximated by a Gaussian distribution.
Reconstruction is most effective when the target image is only a small perturbation from the average illuminating pattern (Fig. 5). This motivates the technique explored in the next section, whereby the target image is modified to leverage this fact.
@@ V.3 Modified target images
If instead we add the average illuminating pattern to the target image before the ghost-projection process is performed, as shown in Fig. 6(a), the NNLS need only focus on producing the features of the modified target image that deviate from this average profile. Many more of the illuminating patterns will, by design, have a greater degree of similarity with this modified image. In this case approximately 3% (or $\sim 1200$) of the patterns are assigned a weight larger than 1% of the maximum weight compared with only 158 with the unmodified target image. Accordingly, Fig. 6(b) shows the improved NNLS ghost-projection construction of this modified target image compared with that for the original target image in Fig. 4(c).
Figure 6: (a) Modified target image, of the capital letter“R” plus the average mask, (b) NNLS construction of modified target image.
The ratio between the intensity of the desired image and the intensity of the average illuminating pattern, in the modified target image, is an adjustable parameter $\beta$. We define $\beta$ such that the image we intend to project is $\beta S+\overline{I_{k}}$ where $S$ is the true desired image. Extreme values of $\beta$ do not produce optimum results: when $\beta$ is large, there is little contribution from the average illuminating pattern, which has been shown to be less successful (this is quantified in Fig. 7), and if $\beta=0$ the desired image is not present at all. This indicates that there is some ideal $\beta$ (although it will depend on the reconstruction quality metric and likely the target image). The modified image used in Fig. 6 has the maximum intensity of the desired image as 50% of the maximum intensity of the average illuminating pattern, as this produced visually successful results.
@@
Figure 7: A comparison between NNLS applied to the original target image and NNLS applied to the modified target image, where the average illuminating pattern has been subtracted from the result. The values in the bottom right corners of each reconstruction are a measure of the residual between the desired image and the reconstruction, minimized with respect to uniform intensity offsets. If the average pattern is “subtracted” by adding $\max\{\overline{I_{k}}\}-\overline{I_{k}}$, then the percentage above the pedestal at which the target image sits is (b) 34%, (d) 27%, and (f) 30%. For a video showing the evolution of frames (c) and (d), see the Supplemental Material [30].
@@
Figure 7 displays the improvement in the reconstruction when the average pattern is added prior to performing NNLS, with the pedestal flattened after reconstruction (Sec. VII details potential methods of flattening the pedestal). Three different $100\times 100$ original target images were chosen for demonstration: (i) a capital letter “R” to test the ghost-projection reconstruction of line-like features (Figs. 7(a-b)), (ii) the first 3 iterations of the Sierpiński carpet fractal to test fine detail and multiple length scales (Figs. 7(c-d) and associated video in Supplemental Material [30]), and (iii) the Southern Cross stellar constellation to test isolated spots akin to the larger spots in the illuminating patterns (Figs. 7(e-f)).
@@
The dimensionless residuals shown in Fig. 7 are defined as the norm squared of the difference divided by the norm squared of the target image. For assessing synthesis quality, this metric is crude in the sense that it is not sensitive to fine detail. For instance, the residual for (c) suggests it is the most faithful of the NNLS reconstructions performed on the original target image, yet it can easily be argued that (a) and (e) capture the majority of the information of their target images, while any details beyond the first iteration of the Sierpiński carpet are lost in (c).
@@
The reconstructions (b), (d), and (f) in Fig. 7 all show significant improvement, both according to the reduced residual and reconstruction of fine detail. Original target images that are initially closer in intensity profile to the pedestal (e.g. (a) and (c)) are recreated more successfully, and adding the average illuminating pattern to the target image has a more pronounced impact for original target images which differ more from the average pattern (e.g. (e) compared with (f)). Additionally, (f) is the most faithfully reconstructed, suggesting that NNLS on the modified image is more successful for images with features that are similar to the illuminating pattern features.
@@
Regarding fine detail, (d) shows that when using the modified target image, the high spatial frequency portions of the diffraction catastrophes (single-pixel feature sizes and smaller) allow for single-pixel information in the target image to be accurately resolved without a substantial increase in the pedestal. Moreover, for all target images tested in this study, when NNLS is applied to the modified target image, the additional pedestal (on top of the average pattern) is spatially constant, so subtracting the average pattern produces a uniform flat background in (b), (d), and (f).
@@
If ghost projection is used to encode data to be read digitally, then the pedestal-subtraction process can be conducted digitally after scanning. Conversely, if a flat pedestal is necessary (e.g. for lithography), then the average pattern needs to be subtracted physically (up to a constant shift). In the latter case, it may be possible to subtract the pedestal during the projection process using an attenuating mask. Alternatively, a suitable intensity distribution may be superimposed afterwards to flatten the pedestal. These ideas are explored in more detail in Sec VII.

### v2

@@ IV.1 Truncated Zernike basis for phase modulations
The beam in Fig. 1(a) is taken to be transversely confined by a circular aperture. Zernike polynomials [8, 25, 26, 27, 28] form a complete orthogonal set over such a region. Additionally, the low-order Zernike polynomials have low spatial frequency. They therefore provide a convenient basis for describing the slowly-varying phase modulations that can be used to generate diffraction catastrophes for the purposes of ghost projection. Working in plane polar coordinates $(\rho,\theta)$, denote the first 21 Zernike polynomials by
@@
To bound the phase modulations, each $c^{k}_{j}$ is randomly chosen from the interval $\tau,\tau]$ for some positive real value $\tau$. In order for the focal-field intensity to contain a caustic network, the value for $\tau$ needs to be sufficiently large. We set the amplitude of the modulated wavefunction to be uniformly 1 within the unit circle and zero elsewhere, thereby simulating a circular aperture that is uniformly illuminated by normally-incident strictly monochromatic plane waves, so
@@
For most of the results in this paper (from Sec. V.3 onwards), a set of $\mathcal{N}=7500$ randomly-generated illuminating patterns was used to construct $100\times 100$ pixel target images. Naturally, $\mathcal{N}=7500$ patterns cannot span the 10 00010\,000-dimensional space of $100\times 100$ pixel images (especially when restricted to only using non-negative coefficients). It is worth noting that this means the set of patterns does not strictly form a basis in the mathematical sense (a linearly independent spanning set). Nonetheless, with the target image modifications explored in Sec. V.3, a set of 7500 patterns is ample to produce successful reconstructions, emphasizing the effectiveness of this technique. In practice, assuming there are no constraints on computation, larger values of $\mathcal{N}$ lead to more successful reconstructions. If $\mathcal{N}$ is far greater than the number of pixels, most vectors will be able to be constructed by numerous different non-negative linear combinations of the illuminating patterns, although these linear combinations may differ in terms of the magnitude of the pedestal they produce. Consequently, as an area of further exploration, it makes sense to talk about choosing the linear combination which maximizes the signal-to-pedestal ratio (which leads to an improved signal-to-noise ratio in the presence of photon shot noise). In this paper, no mechanism for this has been explored.
For $k\in\{1,\cdots,\mathcal{N}\}$ and $j\in\{2,\cdots,21\}$, each of the $c^{k}_{j}$ was uniformly randomly selected from $\tau,\tau]$. The size of $\tau$ was found to be approximately proportional to the average size of the corresponding focal-field illuminating pattern, which is natural on account of the fact that the mean aperture-plane phase gradient is linearly proportional to $\tau$. For this paper, we choose $\tau=9$, as for $100\times 100$ pixel images, this makes the patterns a suitable size for the target images to be synthesized.
@@ IV.2 Focusing the phase modulations via the Fourier transform
A coherent two-dimensional radiation distribution, when propagated infinitely far, becomes the Fourier transform of the initial distribution. A focusing lens enables this to occur at a finite propagation distance, and the Fourier transform appears at the focal plane [8]. In particular, once the modulated wavefunction, $\psi_{k}$, is focused via a lens, the intensity distribution, $I_{k}$, in the focal plane is given by
@@
Here, $\mathcal{F}$ denotes Fourier transformation with respect to transverse Cartesian coordinates $\rho\cos\theta$ and $\rho\sin\theta$, and $A(\rho)$ is an aperture function in scaled (dimensionless) spatial coordinates, defined to be unity for $\rho\leq 1$ and zero elsewhere. The resulting $I_{k}$ are diffraction catastrophes [9, 10, 7]. Here, $\{I_{k}\}_{k\in\{1,\cdots,\mathcal{N}\}}$ serves as the caustic-laden “catastrophe basis” set of illuminating patterns to be used for ghost projection. The process described above is depicted in Fig. 1(a) and some example patterns are shown in Fig. 1(b). For simulation purposes, it is helpful to rasterize the wavefunctions, treating $A(\rho)\exp[i\phi_{k}(\rho,\theta)]$ as a vector with a complex-valued entry for each pixel. This allows the Fourier transform step to be performed computationally as a 2D fast Fourier transform (FFT) [29]. To avoid wrapping of the FFT at the image boundaries, the phase function must be zero-padded [29], i.e., the phase function is given a sufficiently broad border of value zero.
@@ IV.3 Simulating Poisson noise in the focal fields
The number of photons (or electrons, neutrons, etc.) incident on a pixel of the projection plane obeys a Poisson probability distribution [30, 31]. For simplicity, assume that the projecting beam is composed of photons. For the $j$th pixel of the ghost projection, let $\Lambda_{j}:=c\sum_{k}w_{k}[I_{k}]_{j}$ represent the expected number of photons incident on that pixel for some scaling factor $c$. When measured, the number of photons incident at each pixel, $j$, will be Poisson distributed, with mean $\Lambda_{j}$ and standard deviation $\sqrt{\Lambda_{j}}$. To quantify noise levels, a pixel-wise signal-to-noise ratio (SNRj) is introduced, given by
$\text{SNR}_{j}:=\frac{|\Lambda_{j}-P|}{\sqrt{\Lambda_{j}}}$ (19)
@@
$\text{SNR}:=\sqrt{\sum_{j=0}^{N^{2}}\frac{\text{SNR}_{j}^{2}}{N^{2}}}=\frac{1}{N}\left\lVert\frac{\Lambda-P}{\sqrt{\Lambda}}\right\rVert_{2}.$ (20)
@@ V Results and Analysis
In this section we first compare ghost projection performance using both (i) the analytical dot product method, and (ii) the iterative non-negative least-squares method, to determine pattern weights in Sec. V.1. Unlike a random speckle-pattern basis that produces a flat (i.e. uniform) pedestal in target-image synthesis, we observe that a diffraction-catastrophe basis produces a non-uniform pedestal. This observation is presented in Sec. V.2. We then present a method to modify the target image to compensate for this non-uniform pedestal in Sec. V.3. Finally, we explore the effects of Poisson noise on target image quality in Sec. V.4.
V.1 Synthesis of unmodified target images
Figure 3(a) shows an example of a phase function, $\phi_{k}$, that is generated randomly using Zernike polynomials from $\mathcal{Z}$, with the corresponding diffraction catastrophe, $I_{k}$, after focusing shown in Figs. 3(b-c). The spatial scale is in arbitrary units, for all panels. The caustic diffraction catastrophes [7] are evident as the bright lines in the intensity images. While these lines would correspond to nonphysical lines of infinite intensity in the geometric-optics limit of zero wavelength, the finite wavelength softens them to peaked intensities decorated with fine grid-like structures associated with a lattice of alternating phase vortices and antivortices [10].
@@ V.1 Synthesis of unmodified target images
Using a set of $\mathcal{N}=40\,000$ randomly generated aberrated focal-field diffraction-catastrophe patterns as a random basis, the simulated ghost projections in Fig. 4 are produced. The target image, here, is the binary capital letter “R” in Fig. 4(a). Two different ghost-projection algorithms are explored below, namely dot product ghost projection [1] and NNLS ghost projection [2, 3, 4], as described in the following paragraphs (see, also, Sec. III). The dot product reconstruction in Fig. 4(b) has comparatively far lower contrast, with the white section of the target image sitting only approximately 5% above the local pedestal, whereas in the NNLS construction of Fig. 4(c), the white section sits approximately 34% above the local pedestal. This greater contrast is particularly valuable when noise is present (as explored further in Sec. IV.3).
@@
Figure 4: (a) A $100\times 100$ pixel image of the letter “R” used as a target image, (b) a simulated ghost-projection reconstruction of the target image using 40 000 patterns, weighting the catastrophes by their dot product with the target image, (c) a reconstruction using the non-negative least squares algorithm to generate weights, using 40 000 illuminating patterns. The reconstructions have been normalized, so that the minimum value is rendered black and the maximum value is rendered white. (d) There is low correlation between the two methods for generating the weights, $w_{k}$, with the Pearson correlation coefficient $r$ being 0.05.
@@
The “dot product” ghost-projection construction, which follows Ref. [1] by using weights defined in Eq. (11), discards illuminating patterns that have a below-average dot product with the target image. If one thinks of the random illumination patterns as randomly directed vectors in a function space, this method for ghost projection has the simple geometric interpretation that any function-space vector may be synthesized by taking a sheaf of random vectors and retaining only those having a positive projection along the direction of the required vector; the sum of these retained vectors is then approximately parallel to the desired vector (target ghost-projection image). [^6] This geometric construction is illustrated in Fig. 11 of Ref. [1]. Illuminating patterns with lower dot products with the desired image increase the pedestal while having a reduced contribution towards improving the ghost-projection image. Hence the weights of the included illuminating patterns are the dot product between the image and illuminating pattern vector, minus the average dot product. While conceptually clear, the very low contrast of the “dot product” ghost-projection method (see Fig. 4(b)) argues against its utility in practice.
@@
Further comparison of the two ghost-projection approaches may be obtained by plotting each weight, $w_{k}$, obtained using the NNLS approach, versus the corresponding weight obtained via the dot product approach. The resulting scatter plot in Fig. 4(d) reveals a notable disparity between the two reconstruction algorithms, with no significant correlation between the weights $w_{k}$ generated by each procedure. While Pearson’s correlation coefficient ($r=0.05$, stated in Fig. 4(d)) is positive, the support for a linear relationship between the dot product and the NNLS weights is weak, so the dot product cannot easily be used to filter illuminating patterns before the NNLS algorithm is applied. It is worth noting that the dot-product weights are not symmetrically distributed around their mean and are positively skewed. Additionally, the NNLS algorithm only assigns 158 illuminating patterns a weight that is over 1% of the maximum weight (only 0.395% of the 40 000 illuminating patterns), with 62 of these having a dot product below the mean correlation (red) and 96 having a dot product above the mean correlation (blue). This small selection of contributing illuminating patterns is the reason that the NNLS image is noticeably more grainy than the dot product reconstruction, without a large quantity of similarly-weighted illuminating patterns to smooth the image.
@@ V.2 Non-uniform pedestal
Another significant aspect of the images formed in Fig. 4 is the additional background signal, or pedestal; in particular, that it is not spatially uniform. For the case of the dot product reconstruction in Fig. 4(b), the pedestal is approximately the average illuminating pattern (see Fig. 5), while the NNLS construction in Fig. 4(c) has a flatter pedestal of similar profile (as there are far fewer contributing illuminating patterns, the resulting pedestal when using NNLS is more dependent on the target image).
Figure 5: The average diffraction catastrophe, plotted as (a) a 1D profile versus radial distance and (b) a 2D grayscale image. The 2D image above has not been normalized. As may be expected from the central limit theorem, the average illuminating pattern is well approximated by a Gaussian distribution.
Reconstruction is most effective when the target image is only a small perturbation away from the average illuminating pattern (AIP) in Fig. 5. If the mean-corrected target image exhibits only small deviations from zero (i.e. it is relatively flat), adding the AIP to the target image produces a modified target image which is only a small perturbation away from the AIP, as desired. Observing Eq. (14), adding the AIP to the target image achieves a similar effect to replacing the patterns $I_{k}(\mathbf{r})$ in Eq. (13) with $I_{k}(\mathbf{r})-\langle I(\mathbf{r})\rangle$. This is a more sophisticated version of the mean correction in Eq. (14), with the patterns shifted by the AIP rather than a constant value.
This mean correction approach is suboptimal when noise is present (since it produces large pedestals) or when the target image is not relatively flat. In the next section, we will explore a more general tunable technique to account for these scenarios.
@@ V.3 Modified target images
As previous mentioned, the target image $S(\mathbf{r})$ is reconstructed most successfully when its overall spatial profile is similar to $\langle I(\mathbf{r})\rangle$, the AIP. Intuitively, this is because many more patterns will already share large-scale features with the target image, allowing the NNLS to focus on small deviations and prioritize reconstruction of finer details.
To exploit this fact, we propose a simple method for modifying the target image to be more similar to the AIP. Specifically, instead of projecting the target image $S(\mathbf{r})$, we project
$S_{\beta}(\mathbf{r}):=\beta S(\mathbf{r})+(1-\beta)\langle I(\mathbf{r})\rangle,$ (21)
where $\beta\in(0,1]$ is a tunable parameter. Note that $S_{1}(\mathbf{r})=S(\mathbf{r})$. An example of this has been presented in Fig. 6 for $\beta=0.1$. Naturally, when $\beta\neq 1$, it is necessary to “subtract” $(1-\beta)\langle I(\mathbf{r})\rangle$, referred to as the “pedestal”, afterwards to arrive at the desired final image. Methods for this subtraction (or “pedestal flattening”) are explored in Sec. VII.
Figure 6: (a) Modified target image of the capital letter “R” using $\beta=0.1$ in Eq. (21), (b) NNLS reconstruction of this modified target image using 7500 illuminating patterns, (c) reconstruction after the pedestal, namely $(1-\beta)\langle I(\mathbf{r})\rangle$, has been subtracted.
The extreme values for $\beta$ do not lead to optimal results: $\beta=0$ does not have any contribution from the target image, while $\beta=1$ corresponds to the unmodified target image, which is suboptimal in most scenarios (as illustrated in Figs. 7 and 8). To quantify and compare reconstruction quality, we define “dimensionless residuals” as the norm squared of the reconstruction (after pedestal subtraction) minus the target image divided by the norm squared of the target image.
It is possible to optimize the value of $\beta$ by allowing the NNLS algorithm to adjust it as one of its parameters (this typically results in a low-contrast image reconstruction with $\beta\sim 10^{-2}$; the target image is reduced to a slight perturbation about the AIP). To achieve this, the negative of the AIP can be included with the illuminating patterns when the NNLS is conducted, with $\beta$ simultaneously optimized along with the set of exposure times for each pattern (equivalent to setting $\beta=w_{\mathcal{N}+1}$ and $\langle I(\mathbf{r})\rangle=I_{\mathcal{N}+1}$ in Eq. (14)). Although this leads to a mathematically optimal choice of $\beta$ (according to the Euclidean norm), small values of $\beta$ come at the cost of a lower signal-to-pedestal ratio, making the reconstruction more vulnerable to noise in an experimental realization of ghost projection. As a result, choosing $\beta$ is a matter of balancing artifacts with noise. Figure 8 demonstrates an example of this compromise in the presence of Poisson noise, where low values of $\beta$ lead to poor dimensionless residuals due to having low signal-to-noise ratio, while large values of $\beta$ lead to unwanted reconstruction artifacts, resulting in $\beta\approx 0.1$ having a minimal residual. The optimum $\beta$ for any given application will depend on the target image, the noise sources involved, and the chosen reconstruction quality metric (dimensionless residuals as we have defined them may not be the most relevant metric for every application).
Figure 7: NNLS reconstructions of a Sierpiński carpet using different $\beta$ values (see Eq. (21)). Dimensionless residuals are included in the bottom right corner of each image. To simulate Poisson noise (photon shot noise), the projection’s exposure is such that, on average, a maximum-exposure pixel of the target image receives 1000 photons more than a minimum-exposure pixel at the same location. The total number of photons hitting a pixel at position $\mathbf{r}$ is $N(\mathbf{r})=\frac{1000}{\beta}(1-\beta)\langle I(\mathbf{r})\rangle+1000S(\mathbf{r})$. Since the standard deviation of the Poisson noise is $\sqrt{N(\mathbf{r})}$, the SNR is lower for smaller $\beta$. There is thus a compromise between SNR and reconstruction quality, as illustrated in the third column.
Figure 8: Dimensionless residuals of reconstructions of a Sierpiński carpet as a function of $\beta$. In the absence of noise (blue), optimum values are typically $\beta\sim 10^{-2}$. However, Poisson noise (photon shot noise) is more impactful for small $\beta$. To simulate the Poisson noise (orange), the projection exposure is set such that a maximum-exposure pixel (“white pixel”) of the original target image receives an average of 1000 photons more than a minimum-exposure pixel at the same location. This leads to an optimum of $\beta\approx 0.1$.
A further consideration when choosing $\beta$ for practical applications is that the number of patterns given a significant weight by the NNLS also depends on the value of $\beta$. Smaller $\beta$ values tend to result in many patterns being given similar weights, while larger values of $\beta$ result in only a small number of patterns being given large weights. Recall, in this context, that weights typically correspond to exposure times. Hence, smaller values of $\beta$ require more precise exposure times in practical implementations, while larger values of $\beta$ lead to experiments that are more robust to uncertainty in exposure times.
@@
Figure 9: A comparison between NNLS applied to the original target image (i.e. $\beta=1$) and NNLS applied to the modified target image (with $\beta=0.1$), where the AIP has been subtracted from the result. The values in the bottom right corners of each reconstruction are a measure of the residual between the desired image and the reconstruction, minimized with respect to uniform intensity offsets. For a video showing the ghost projection process for panel (d), see the Supplemental Material [32].
@@
Figure 9 displays the improvement in reconstruction fidelity when various target images are modified according to Eq. (21), using $\beta=0.1$. The images are constructed using NNLS, and for $\beta=0.1$, the pedestals are flattened after reconstruction (Sec. VII details potential practical methods for flattening the pedestal). Three different $100\times 100$ original target images were chosen for demonstration: (i) a capital letter “R” to test the ghost-projection reconstruction of line-like features (Figs. 9(a-b)), (ii) the first 3 iterations of the Sierpiński carpet fractal to test fine detail and multiple length scales (Figs. 9(c-d) and associated video in Supplemental Material [32]), and (iii) the Southern Cross stellar constellation to test isolated spots akin to the larger spots in the illuminating patterns (Figs. 9(e-f)).
@@
Recall that the dimensionless residuals shown in Fig. 9 are defined as the norm squared of the difference divided by the norm squared of the target image. For assessing synthesis quality, this metric is crude in the sense that it is not sensitive to fine detail. For instance, the residual for (c) suggests it is the most faithful of the NNLS reconstructions performed on the original target image, yet it can easily be argued that (a) and (e) capture the majority of the information of their target images, while any details beyond the first iteration of the Sierpiński carpet are lost in (c).
@@
The reconstructions (b), (d), and (f) in Fig. 9 all show significant improvement, both according to the reduced residual and reconstruction of fine detail. Original target images that are initially closer in intensity profile to the pedestal (e.g. (a) and (c)) are recreated more successfully, and adding the AIP to the target image has a more pronounced impact for original target images which differ more from the average pattern (e.g. (e) compared with (f)).
@@
Regarding fine detail, Fig. 9(d) shows that when using the modified target image, the high spatial frequency portions of the diffraction catastrophes (single-pixel feature sizes and smaller) allow single-pixel information in the target image to be accurately resolved. Moreover, for all target images tested in this study, when NNLS is applied to the modified target image, the additional pedestal (on top of the AIP) is spatially constant, so subtracting the AIP produces a uniform flat background in Figs. 9(b), (d), and (f).
@@
If ghost projection is used to encode data to be read digitally, then the pedestal-subtraction process can be conducted digitally after scanning. However, if a flat pedestal is necessary (e.g. for lithography), then the AIP needs to be subtracted physically (up to a constant shift). In the latter case, it may be possible to subtract the pedestal during the projection process using an attenuating mask. Alternatively, a suitable intensity distribution may be superimposed afterwards to flatten the pedestal. These ideas are explored in more detail in Sec. VII.